OpenStudy (anonymous):

what numbers have a product of 9/16 and a sum of -3/2?

undeadknight26 (undeadknight26):

Are there choices?

OpenStudy (anonymous):

no

undeadknight26 (undeadknight26):

@SolomonZelman

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \sf a \times b=9/16 }\) \(\Large\color{blue}{ \sf a + b=-3/2 }\) if this is because you need to solve a quadratic equation, then use a quadratic formula instead, don't factor (unless you really have to)

OpenStudy (solomonzelman):

can you solve for a and b ? (doesn't matter which one is which)

OpenStudy (anonymous):

i dont understand

OpenStudy (solomonzelman):

You don't why I wrote what I wrote in blue ?

OpenStudy (anonymous):

well i mean its just that, how do u solve for a and b and there isnt two integers

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{9}{16} =a \times b}\) \(\Huge\color{blue}{ \sf -\frac{3}{2} =a + b}\) red is the equation that results after I divide the first blue equation by b on both sides.\(\Huge\color{red}{ \sf \frac{9}{16b} =a }\)

OpenStudy (solomonzelman):

I got disconnected...

OpenStudy (solomonzelman):

Substitute the left side of the red equation above, for a. \(\Huge\color{red}{ \sf -\frac{3}{2}= \frac{9}{16b}+b }\) \(\Huge\color{red}{ \sf -\frac{3}{2}= \frac{9}{16b}+\frac{16b^2}{16b} }\) \(\Huge\color{red}{ \sf -\frac{3}{2}= \frac{9+16b^2}{16b} }\)

OpenStudy (solomonzelman):

Cross multiply. \(\Huge\color{red}{ \sf -\frac{3}{2}= \frac{9+16b^2}{16b} }\) \(\Huge\color{red}{ \sf \frac{-3}{2}= \frac{9+16b^2}{16b} }\) \(\Huge\color{red}{ \sf -48b=18+36b^2}\)

OpenStudy (solomonzelman):

\(\Huge\color{red}{ \sf -48b=18+36b^2}\) \(\Huge\color{red}{ \sf 36b^2-48b+18=0}\) \(\Huge\color{red}{ \sf 18b^2-24b+9=0}\) can you solve this quadratic ?

OpenStudy (anonymous):

the answer is 3/4

OpenStudy (anonymous):

thats what the book told me but i dont know how they got it